OpenStudy (anonymous):

solve (2sinᶱ-1= 3/sinᶱ) using u-substitution

OpenStudy (anonymous):

u = sinx. Then we obtain: \[2u - 1 = \frac {3}{u}\] With \[u \neq 0\], we can multiplied both sides by u \[2u^2 - u - 3 = 0 (u \neq 0)\] Solve it and you have \[u_1 = -1 ; u_2 = \frac{3}{2}\] Substitute into sinx, you can calculate x by yourself.